OpenStudy (marziajackson$):

What is the value of the expression 2/4^-3 ? A. 1/32 B. 1/6 C. 128 D. 256

OpenStudy (michele_laino):

hint: \[\Large \frac{2}{{{4^{ - 3}}}} = 2 \cdot {4^3} = ...?\]

OpenStudy (marziajackson$):

is it A?

OpenStudy (michele_laino):

hint: \[\Large \frac{2}{{{4^{ - 3}}}} = 2 \cdot {4^3} = 2 \cdot 64 = ...?\] since 4^3=4*4*4=64

OpenStudy (marziajackson$):

Is the answer C?

OpenStudy (michele_laino):

correct!

OpenStudy (marziajackson$):

@Alicia_redd